# MINGLE: VLMs for Semantically Complex Region Detection in Urban Scenes

### v2

@@ 5.1 Fine-tuning
Our experiments with full parameter fine-tuning showed promising results; however, given the relatively small training dataset size and resource limitations, they are currently excluded from the scope of this study.
5.2 Social Group Region Detection
@@
Table 3: Currently available VLMs are unable to reliably detect semantically complex regions
5.3 Distance and Depth Filtering Effects
@@
Table 4: Summary of depth and distance filtering effects on the social group region detection
@@
The results demonstrate that the optimal effectiveness is attained at distance values 0.3–0.5 and depth values 80–100. However, the limited size of the benchmark dataset may limit the generalizability of this finding and we hence recommend re-running a similar analysis on a relevant dataset of interest.
5.4 Social Group Region Dataset
@@
6 Conclusion
In this work, we addressed the previously unsolved urban analysis challenge: social group region detection. Our extensive zero-shot evaluation confirmed that existing models fail to reliably localize social groups, highlighting the semantic complexity of the task. To tackle this, we proposed MINGLE: a modular three-stage VLM-based pipeline, which reframes the problem from direct region-level detection to pairwise social affiliation classification.
To encourage future work in both urban planning and semantically complex region detection, we release a large-scale street-view dataset annotated with granular information about social groups created through combination of manual and pipeline-based labeling.

### v3

@@ Results
Our experiments with full parameter fine-tuning showed promising results. However, given the relatively small training dataset size and resource limitations, they are currently excluded from the scope of this study.
Social Group Region Detection
@@
Table 3: Currently available VLMs are unable to reliably detect semantically complex regions.
Distance and Depth Filtering Effects
@@
Table 4: Summary of depth and distance filtering effects on the social group region detection.
@@
The results demonstrate that the optimal effectiveness is attained at distance values 0.3–0.5 and depth values 80–100. However, the limited size of the benchmark dataset may limit the generalization of this finding, and we recommend re-running a similar analysis on a relevant dataset of interest.
Social Group Region Dataset
@@
Conclusion
In this work, we address the previously unsolved urban analysis challenge: social group region detection. Our extensive zero-shot evaluation confirms that existing models fail to reliably localize social groups, highlighting the semantic complexity of the task. To tackle this, we have presented MINGLE: a modular three-stage VLM-based pipeline, which reframes the problem from direct region-level detection to pairwise social affiliation classification.
To encourage future work, we release a large-scale street-view dataset annotated with granular information about social groups created through a combination of manual and pipeline-based labeling.
Acknowledgments
The project was supported by the Designing for Sustainability research program, a joint initiative of the Morningside Academy for Design at MIT and the Hasso Plattner Institute, generously funded by the Hasso Plattner Foundation.
This work used the Jetstream2 GPU and NCSA Delta GPU resources through allocation CIS250088 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, supported by the U.S. National Science Foundation (NSF) under grants #2138259, #2138286, #2138307, #2137603, and #2138296.